A city is in the process of expanding their water system. They just built a large water tower with a height of 100m about the city water mains. What size of water main pipe should they use so that there is at least 400 kPA or pressure in the pipe when 4m^3/s of water is flowing through the pipe? The density of water is 1,000 kg/m^3.
Answer #1

let's denote the density as d.

d=1000 kg/m^3

Applying Bernoulli's equation,

P1 + dgh= P2 + 1/2*d*v^2, P1=100kpa. P2=400 kpa,h=100,

g=9.8

v=36.8m/s

let's assume radius of pipe is r.So,

pi*r^2 * v= 4, Mass conservation

r=18.6 cm